# MorphoSeg: An Uncertainty-Aware Deep Learning Method for Biomedical Segmentation of Complex Cellular Morphologies

### v1

@@ 3.3 Unknown-aware Training Objective
$\mathcal{L}_{\text{Dice\_out}}(f_{\text{syn}}(\mathbf{x};\theta),y)$ $=1-\frac{2\sum_{i}p_{\text{syn},i}g_{i}+\epsilon}{\sum_{i}p_{\text{syn},i}^{2}+\sum_{i}g_{i}^{2}+\epsilon},$ (13)
@@
where $p_{\text{syn},i}$ represents the predicted probabilities for the actual class labels $g_{i}$ from the synthetic outputs, and $f_{\text{out}}(\mathbf{x}_{i};\theta)$ denotes the model’s output for synthetic data designed to simulate under-represented features.
@@ 4.1 Data Description
The NT2 dataset consists of 105 high-resolution raw images of NT2 cells undergoing differentiation, of which 36 images are fully annotated. These images have resolutions primarily around $4000\times 3000$ pixels, enabling detailed morphological analysis and accurate segmentation.
@@ 4.2 Data Preprocessing
Data Augmentation. To efficiently train our TransUNet model, we have devised a method to augment the dataset by transforming the high-resolution images into smaller, manageable patches. This is achieved by dividing the images into patches of sizes $224\times 224$, $448\times 448$, $1000\times 1000$, $1500\times 1500$, and $2000\times 2000$ pixels. We employ an overlapping strategy with an overlap ratio of 0.35 to preserve the spatial context across adjacent patches, which is crucial for capturing the comprehensive spatial information of the cell structures. Combined with the original undivided images and masks, this makes the full dataset for training and testing.
Using this technique, we generate a total of 15,710 patches. These patches are then randomly split into a training set and a testing set, maintaining an 80% to 20% ratio, respectively. This distribution results in 12,568 patches for training and 3,142 patches for testing. To optimize the model’s performance, we implement a data pruning strategy during both the training and testing stages, excluding all patches that do not contain any part of the annotated masks, which will reduce the training set to 10,016 samples for testing.
@@ 5.1 Implementation details
We evaluate our model against established benchmarks including Cellpose and TransUNet, utilizing the AdamW optimizer with exponential learning rate decay on the R50-ViT-B architecture. Specifically for Cellpose, we utilize two of the highest performing pre-trained configurations, cyto3 and cyto2, both set with a cell diameter of $30$. These models are trained with a learning rate of $0.01$, a weight decay of $5\times 10^{-5}$, and over $200$ epochs. The batch size for training is set at $8$, with verbose output enabled for detailed progress tracking. During testing, we adapted the Cellpose model to handle larger cells by setting the diameter parameter to $100$, with no flow threshold and a cell probability threshold of $0.2$, adjusting the channels accordingly to match specific imaging conditions. For the TransUNet, we adhere which includes two types of encoders: a pure Transformer-based encoder, denoted as “ViT-B16,” and a hybrid encoder that combines ResNet-50 with ViT, denoted as “R50-B16”. The input resolution and patch size $P$ are set to 224$\times$224 and 16, respectively. We set the learning rate to 0.01, momentum to 0.9, and weight decay to $1\times 10^{-4}$. The models are trained over 200 epochs, with sampling commencing at the 150th epoch. Both models are trained on two NVIDIA A100 GPUs with a batch size of 256, focusing on 10,000 critical samples by selectively sampling 1,000 pixels per image to enhance training efficiency and effectiveness.
@@ 5.2 Evaluation Metrics
To evaluate our method, we considered four key metrics: 1) the average Dice Similarity Coefficient (DSC) [53], 2) the average Hausdorff Distance (HD) [54], 3) the Intersection over Union (IoU) [55], and 4) the mean Average Precision (mAP) [56]. These metrics provide a comprehensive assessment of segmentation accuracy and model performance.
For each class, the unique matching of $(y,\hat{y})$ splits the predicted and the GT segments into three sets:
  • True Positives (TP): the number of correctly predicted pixels that match the ground truth.
  • False Positives (FP): the number of pixels incorrectly predicted as part of the object when they are not in the ground truth.
  • False negatives (FN): the number of pixels that are missed by the prediction but are present in the ground truth.
The Dice Similarity Coefficient (DSC) is a standard metric for assessing the overlap between the predicted segmentation mask and the ground truth mask. It is defined as:
DSC $=\frac{2\times\text{TP}}{2\times\text{TP}+\text{FP}+\text{FN}},$ (17)
where TP, FP, and FN denote the number of true positive, false positive, and false negative pixels, respectively. The Mean DSC is computed by averaging the DSC values across all segmentation instances or classes in the dataset. The Mean DSC provides a robust measure of overall segmentation accuracy, facilitating a comprehensive evaluation of model performance across all classes or instances. Since we only have $1$ target class which is the cell, we only use the DSC as the evaluation metric.
To further assess the boundary accuracy, we employ the 95th Percentile Hausdorff Distance (HD95), which offers a more robust measure of the spatial discrepancy between the predicted and ground truth segmentation boundaries. The Hausdorff Distance (HD) between two sets, $X$ and $Y$, is given by:
$\text{HD}(X,Y)$ $=\max(h(X,Y),h(Y,X)),$ (18)
where
$h(X,Y)=\max_{x\in X}\min_{y\in Y}d(x,y),$ (19)
and $d(x,y)$ denotes the Euclidean distance between points $x\in X$ and $y\in Y$.
The HD95 is defined as the 95th percentile of the sorted set of all distances between the predicted and ground truth boundaries:
@@
where $d(x,Y)=\min_{y\in Y}d(x,y)$ and $d(y,X)=\min_{x\in X}d(y,x)$. The HD95 metric reduces the influence of outliers by focusing on the 95th percentile distance, providing a robust measure of boundary alignment between the predicted and ground truth segmentations.
The model evaluation should not only use the statistic metrics but also need a qualitative evaluation from the actual segmentation view. The intersection over union (IoU) [57] as an important evaluation index in semantic segmentation measures the overlap of the ground truth and prediction region. It is also calculated:
IoU=Ground TruthPredictionGround $\text{IoU}=\frac{\text{Ground Truth}\bigcap\text{Prediction}}{\text{Ground Truth}\bigcup\text{Prediction}},$ (21)
where $\bigcap$ denotes the intersection operator and $\bigcup$ represents the union operator. The IoU is generally calculated based on categories, which is to accumulate the IoU value of each category. The IoU value is to average the sum IoU results of each category to obtain a global evaluation. Therefore, the IoU is actually the mean value, that is, the average crossover ratio (mean IoU).
Finally, the mean Average Precision (mAP) is defined as:
mAP $=\frac{1}{N}\sum_{i=1}^{N}\text{AP}_{i},$ (22)
where $\text{AP}_{i}$ represents the average precision for the $i$-th class, and $N$ is the total number of classes.
These metrics collectively enable a robust evaluation of our model’s performance in terms of segmentation accuracy and overall effectiveness.
@@ 5.3 Comparison with Baselines
Table. 1 presents the main results of the proposed approach, which includes training with outliers compared to the Cellpose pretrained on cyto3 and cyto2, and TransUNet of ViT-B16 and R50-B16 variants. Despite the advantages of the Cellpose model, such as requiring a smaller number of annotations, it does not offer superior performance due to its reliance on the flow and diameter of cells. In our study, these characteristic properties vary across different stages, and the irregular shapes of cells do not provide consistent diameters for effective segmentation. In comparison, our training approach with Outlier Synthesis enhances the performance significantly over the state-of-the-art TransUNet across all types of loss combinations.
Method DSC HD95 IoU-0.5 IoU-0.75 IoU-0.9 mAP
Framework Encoder Loss $\uparrow$ $\downarrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$
Cellpose [6] cyto3 63.66 137.9 25.31 12.35 3.641 48.51
cyto2 48.15 181.5 19.02 8.220 2.569 36.78
TransUnet [1] ViT-B16 72.47 38.21 76.42 30.19 3.019 36.54
R50-B16 80.35 21.98 88.00 58.67 9.220 51.97
MorphoSeg ViT-B16 Balance 72.78 38.69 77.35 30.19 3.787 37.46
Norm 74.24 36.82 81.54 43.82 4.765 38.96
Pareto 73.67 37.10 80.43 42.66 3.433 37.58
R50-B16 Balance 86.33 16.33 93.02 79.68 20.11 64.27
Norm 84.40 26.37 91.64 72.50 11.30 58.48
Pareto 86.57 15.75 93.02 80.01 21.87 64.96
@@
Table 1: Comparison on our proposed cell dataset. All methods incorporating Outlier Synthesis are trained with a sample size of $100,000$, and the selection count is set to $10,000$. $\uparrow$ indicates that larger values are preferable, while $\downarrow$ indicates that smaller values are better. All numerical values are presented as percentages. Bold numbers indicate superior results.
@@ 5.4 Ablation Studies
Figure 4: Visualization of Ntera-2 Cell Segmentation in the Testing Dataset: Original Images, Segmentation Masks, and Overlays. Row A displays the original images, Row B displays the segmentation mask and Row C displays the mask overlay on the original image. Image set 1 highlights the ability of the network to segment cells in an unfocused image. Image set 2 highlights the ability of the network to segment cells despite low contrast and unclear cell borders. Images 3 and 4 display ability of the network to segment high-quality images at different magnifications, x10 and x25, respectively.

### v2

@@ 3.3 Uncertainty-aware Training Objective
$\mathcal{L}_{\text{Dice\_out}}(f_{\text{out}}(\mathbf{x};\theta),y)$ $=1-\frac{2\sum_{i}\hat{y}_{\text{syn},i}y_{i}+\epsilon}{\sum_{i}\hat{y}_{\text{syn},i}^{2}+\sum_{i}y_{i}^{2}+\epsilon},$ (13)
@@
where $\hat{y}_{\text{syn},i}$ represents the predicted probabilities for the actual class labels $y_{i}$ from the synthetic outputs, and $f_{\text{out}}(\mathbf{x}_{i};\theta)$ denotes the model’s output for synthetic data designed to simulate under-represented features.
@@ 4.1 Data Description
The NT2 dataset consists of 105 high-resolution raw images of NT2 cells undergoing differentiation, of which 36 images are fully annotated, including 2425 individual cells annotated. These images have resolutions primarily around $4000\times 3000$ pixels, enabling detailed morphological analysis and accurate segmentation.
@@ 4.2 Data Preprocessing
Data Augmentation. We randomly partition the entire dataset into two subsets, allocating 80% for training and 20% for testing. To effectively train our TransUNet-based MorphoSeg model, we augment the data by partitioning each high-resolution image into smaller, manageable patches. Specifically, we extract patches of sizes $224\times 224$, $448\times 448$, $1000\times 1000$, $1500\times 1500$, and $2000\times 2000$ pixels using an overlapping strategy with a 35% overlap. This approach preserves the spatial context across adjacent patches, which is essential for capturing the complex spatial relationships inherent in cell structures. In addition to these patches, the original full-resolution images and their corresponding masks are included to form the complete training and testing dataset.
Using this technique, we generate a total of 15,710 patches. This distribution results in 12,568 patches for training and 3,142 patches for testing. To optimize the model’s performance, we implement a data pruning strategy during both the training and testing stages, excluding all patches that do not contain any part of the annotated masks, which will reduce the training set to 10,016 samples for training.
In addition to evaluating performance on the patch-based testing set, we also assess our model on full-resolution images that lack annotations (i.e., the 69 images not included in the 36 annotated images out of 105 total). This evaluation more closely reflects real-world scenarios where researchers directly apply the trained model to raw images, yielding segmentation masks for the entire image. Visualization examples of these results are presented in Figure 4.
@@ 5.1 Implementation details
We evaluate our model against established benchmarks including Cellpose and TransUNet, utilizing the AdamW optimizer with exponential learning rate decay on the R50-ViT-B architecture. Specifically for Cellpose, we utilize two of the highest performing pre-trained configurations, cyto3 and cyto2, both set with a cell diameter of $30$. These models are trained with a learning rate of $0.01$, a weight decay of $5\times 10^{-5}$, and over $200$ epochs. The batch size for training is set at $8$, with verbose output enabled for detailed progress tracking. During testing, we adapted the Cellpose model to handle larger cells by setting the diameter parameter to $100$, with no flow threshold and a cell probability threshold of $0.2$, adjusting the channels accordingly to match specific imaging conditions. For the TransUNet, we adhere which includes two types of encoders: a pure Transformer-based encoder, denoted as “ViT-B16,” and a hybrid encoder that combines ResNet-50 with ViT, denoted as “R50-B16”. The input resolution and patch size $P$ are set to 224$\times$224 and 16, respectively. We set the learning rate to 0.01, momentum to 0.9, and weight decay to $1\times 10^{-4}$. All models are trained over 200 epochs, with sampling commencing at the 150th epoch. Batch sizes of 128 and 256 are implemented depending on the model size, focusing on 10,$0$ critical samples by selectively sampling 1,$0$ pixels per image to enhance training efficiency and effectiveness.
@@ 5.2 Evaluation Metrics
We evaluate our method using four key metrics: the average Dice Similarity Coefficient (DSC) [44] for assessing overlap accuracy, the 95th Percentile Hausdorff Distance (HD95) [45] for measuring boundary alignment, the Intersection over Union (IoU) [46] for quantifying region overlap, and the mean Average Precision (mAP) [47] for overall object detection performance. These metrics, calculated from true positives, false positives, and false negatives, offer a comprehensive view of segmentation quality, while qualitative evaluations further validate the practical effectiveness of our approach.
Method Model DSC $\uparrow$ HD95 $\downarrow$ IoU-0.5 $\uparrow$ IoU-0.75 $\uparrow$ IoU-0.9 $\uparrow$ mAP $\uparrow$
Cellpose [6] cyto3 63.66$\pm$1.23 137.90$\pm$0.95 25.31$\pm$1.11 12.35$\pm$0.87 3.64$\pm$1.45 13.77$\pm$0.67
cyto2 48.15$\pm$0.56 181.50$\pm$1.98 19.02$\pm$1.34 8.22$\pm$0.75 2.57$\pm$1.05 9.94$\pm$0.62
UNet++ [21] - 76.24$\pm$1.29 39.53$\pm$1.70 82.24$\pm$1.12 43.46$\pm$1.05 4.15$\pm$0.60 43.28$\pm$0.55
@@
SwinUnet [25] - 72.78$\pm$1.48 42.77$\pm$1.95 77.29$\pm$1.10 31.36$\pm$0.95 3.27$\pm$0.65 37.31$\pm$0.53
[…]
ScaleFormer [26] - 72.13$\pm$1.38 44.09$\pm$1.90 75.54$\pm$1.10 33.28$\pm$0.95 2.67$\pm$0.60 37.16$\pm$0.52
[…]
TopoSeg [33] ViT-B16 62.08$\pm$1.12 96.51$\pm$1.33 58.30$\pm$0.78 10.93$\pm$0.14 1.28$\pm$0.03 23.50$\pm$0.26
R50-B16 66.14$\pm$1.55 97.77$\pm$1.23 64.21$\pm$1.67 15.12$\pm$0.78 1.83$\pm$1.11 27.05$\pm$0.72
[…]
TransUNet [1] ViT-B16 72.97$\pm$1.50 32.94$\pm$1.20 78.33$\pm$1.11 34.52$\pm$0.87 3.51$\pm$0.89 38.79$\pm$0.56
R50-B16 80.35$\pm$1.67 21.98$\pm$0.99 88.00$\pm$1.11 58.67$\pm$1.45 9.22$\pm$1.02 51.96$\pm$0.70
[…]
Table 1: Comparison on our proposed cell dataset. All methods incorporating outlier synthesis are trained with a sample size of $100,000$, and the selection count is set to $10,000$. $\uparrow$ indicates that larger values are preferable, while $\downarrow$ indicates that smaller values are better. All numerical values are presented as percentages. Bold numbers denote the best performance, with results presented as mean $\pm$ standard deviation over 3 runs.
@@ 5.3 Comparison with Baselines
Table 1 presents a comprehensive comparison of our proposed approach, MorphoSeg, against several baselines including Cellpose, TransUNet, and TopoSeg UNet++ , SwinUNet and ScaleFormer. Cellpose, a CNN-based U-Net model tailored for cell segmentation, is evaluated in its cyto3 and cyto2 variants. Although Cellpose benefits from requiring fewer annotations, its performance is hindered by its reliance on cell flow and diameter estimations—characteristics that vary significantly across different cell stages. This variability is reflected in its DSC scores of $63.66$ (cyto3) and $48.15$ (cyto2), indicating limitations when applied to datasets with diverse cell morphologies.
In addition to Cellpose, we integrated MorphoSeg with another CNN-based architecture, UNet++ [21]. To maintain consistent experimental settings, we sampled the virtual outliers at the output level of the segmentation map rather than at a reduced dimensionality in the latent space. Although sampling in the latent space might simplify training and further improve results, even under these suboptimal conditions, our MorphoSeg still yielded a measurable improvement for UNet++, as demonstrated in Table 1.
Furthermore, we extended MorphoSeg to two state-of-the-art Transformer-based segmentation backbone networks, SwinUNet [25] and ScaleFormer [26]. The results reported in Table 1 were obtained using the official default settings for these models. The integration of MorphoSeg led to significant performance gains across multiple evaluation metrics for both SwinUnet and ScaleFormer, thereby demonstrating the robustness and versatility of our approach.
TransUNet was evaluated in two variants: ViT-B16 and R50-B16. The ViT-B16 variant exhibits suboptimal performance with a DSC of only $25.00$, likely due to challenges in distinguishing cells from the background under varying lighting conditions. In contrast, the R50-B16 variant demonstrates markedly better performance, achieving a DSC of $80.35$. This improvement underscores the benefit of incorporating a ResNet-50-based encoder for effectively capturing low-level features in high-resolution medical images.
We implement TopoSeg based on TransUNet. As a state-of-the-art semi-supervised method that leverages topological consistency, it shows moderate performance. Its ViT-B16 variant attains a DSC of $26.53$, while the R50-B16 variant improves to $66.14$. However, across all metrics, TopoSeg remains less competitive compared to the best-performing models. The configurations used for TopoSeg remain the default settings.
Our proposed method, MorphoSeg, further enhances segmentation accuracy by incorporating virtual outlier sampling and an uncertainty-aware training objective. When combined with the R50-B16 encoder, MorphoSeg achieves a DSC of $86.57$, a HD95 of $15.75$, and superior IoU and mAP scores compared to all baselines. These results highlight the effectiveness of our approach in managing the inherent variability in cell morphology. Additionally, while an integration of our method with TopoSeg was explored using both ViT-B16 and R50-B16 variants, the standalone performance of MorphoSeg with the R50-B16 encoder remains the most competitive, emphasizing its robust capability for accurate cell segmentation.
[…]
MorphoSeg ViT-B16 Balance 76.20$\pm$1.50 31.77$\pm$3.00 65.34$\pm$1.00 39.78$\pm$0.90 4.07$\pm$0.50 36.40$\pm$0.48
Norm 84.89$\pm$1.56 15.37$\pm$1.78 92.50$\pm$1.23 74.02$\pm$1.36 18.08$\pm$0.12 61.53$\pm$0.61
Pareto 79.08$\pm$1.34 26.20$\pm$1.11 69.01$\pm$1.45 49.92$\pm$1.67 5.95$\pm$0.67 41.63$\pm$0.77
R50-B16 Balance 86.33$\pm$1.11 16.33$\pm$0.99 93.02$\pm$1.45 79.68$\pm$1.42 20.11$\pm$1.67 64.27$\pm$0.88
Norm 84.40$\pm$1.16 26.37$\pm$1.11 91.64$\pm$1.22 72.50$\pm$1.67 11.30$\pm$1.34 58.48$\pm$0.82
Pareto 86.57$\pm$1.34 15.75$\pm$1.11 93.02$\pm$2.45 80.01$\pm$1.93 21.87$\pm$0.67 64.97$\pm$1.06
@@
Table 2: Ablation study on the effect of different loss function designs for MorphoSeg. All methods incorporating outlier synthesis are trained using the same configuration as in Table 1. $\uparrow$ indicates that larger values are preferable, while $\downarrow$ indicates that smaller values are better. Bold numbers denote the best performance, with results presented as mean $\pm$ standard deviation over 3 runs.
[…]
ViT-B16 12 12 $\times$ Balance 76.20$\pm$1.50 31.77$\pm$3.00 42.31$\pm$0.69
Norm 84.89$\pm$1.56 15.37$\pm$1.78 61.53$\pm$0.61
Pareto 79.08$\pm$1.34 26.20$\pm$1.11 47.58$\pm$0.55
ViT-L16 24 16 $\times$ Balance 77.35$\pm$1.45 30.12$\pm$2.85 43.27$\pm$0.74
Norm 85.42$\pm$1.59 14.88$\pm$1.77 62.34$\pm$0.68
Pareto 79.97$\pm$1.38 25.83$\pm$1.12 48.21$\pm$0.63
R50-B16 12 12 $\checkmark$ Balance 86.33$\pm$1.11 16.33$\pm$0.99 64.27$\pm$0.77
Norm 84.40$\pm$1.16 26.37$\pm$1.11 58.48$\pm$0.64
Pareto 86.57$\pm$1.34 15.75$\pm$1.11 64.97$\pm$1.06
R50-L32 32 16 $\checkmark$ Balance 88.50$\pm$1.26 14.50$\pm$1.72 67.32$\pm$0.71
Norm 87.55$\pm$1.15 17.82$\pm$1.19 62.75$\pm$0.68
Pareto 88.95$\pm$1.34 14.25$\pm$1.21 68.11$\pm$0.77
Table 3: Ablation studies on different network architectures. For each encoder variant, we report the number of Transformer layers, attention heads, and indicate whether the encoder is hybrid (ResNet backbone; $\checkmark$) or pure Transformer ($\times$). Results are presented for three loss variants: Balance, Norm, and Pareto.
@@ 5.4 Ablation Studies
BloodCell[49] LIVECell[8]
Method Loss DSC $\uparrow$ HD95 $\downarrow$ IoU-0.5 $\uparrow$ IoU-0.75 $\uparrow$ IoU-0.9 $\uparrow$ mAP $\uparrow$ DSC $\uparrow$ HD95 $\downarrow$ IoU-0.5 $\uparrow$ IoU-0.75 $\uparrow$ IoU-0.9 $\uparrow$ mAP $\uparrow$
UNet++[21] - 96.11 2.21 100.00 100.00 83.02 94.34 88.52 7.55 100.00 71.93 15.79 62.57
[…]
TransUNet [1] - 90.56 4.80 100.00 98.74 0.00 66.25 73.83 18.08 77.19 19.30 0.00 32.16
[…]
SwinUNet [25] - 95.37 2.87 100.00 99.37 69.81 89.73 82.95 9.37 100.00 40.35 3.51 47.95
[…]
ScaleFormer [26] - 95.91 2.17 100.00 100.00 78.62 92.87 88.48 6.52 100.00 71.93 12.28 61.40
[…]
Table 4: Segmentation results on the BloodCell and LIVECell datasets of several benchmarks. The results of using MorphoSeg are reported with three loss functions: Balance, Norm, and Pareto. The TransUNet presented in this table uses the hybrid encoder of R50-ViT-B-16. Best results for each dataset for each method is marked in Bold numbers.
Table 2 shows an ablation study comparing three loss strategies for MorphoSeg using two encoder backbones, ViT-B16 and R50-B16. With ViT-B16, the Balance loss yields slightly higher DSC and lower HD95 while all methods register low IoU values at high thresholds. In contrast, the R50-B16 encoder achieves much better performance overall and the Pareto strategy delivers the best results with the highest DSC and mAP and the lowest HD95. These results suggest that the choice of loss is more influential with R50-B16 and that the Pareto approach offers a modest advantage.
Figure 4: Visualization of full input images with model outputs. The presented images are the direct outputs from the model without annotations or ground truth masks, representing unseen data during training and testing. Due to the high resolution, differences among the outputs of TransUNet, TopoSemi, and MorphoSeg are not easily discernible. For a more detailed comparison, zoomed-in results are provided in Figure 5.
Ablation Studies on Different Network Architectures. To evaluate the effect of different network components, we performed experiments on various architectures with differing numbers of Transformer layers and attention heads. Table 3 summarizes these results, which show that larger networks generally yield better performance. Specifically, the “R50-B16” model represents our hybrid architecture combining a ResNet-50 backbone with a Vision Transformer configured in a base setting, consisting of 12 Transformer layers with 12 self-attention heads per layer and a patch size of 16. In contrast, the “R50-L32” model uses the same ResNet-50 backbone but incorporates a larger Transformer module with 32 layers and 16 self-attention heads per layer. For the pure Vision Transformer architectures, ViT-B16 comprises 12 layers with 12 attention heads per layer, while ViT-L16 features 24 layers with 16 attention heads per layer.
5.5 Experiments on Additional Datasets
To further validate the generalization and robustness of our proposed MorphoSeg, we performed experiments on two widely recognized datasets: the BloodCell dataset [49] and the LIVECell dataset [8]. The BloodCell dataset (BBBC041v1) consists of 1,364 microscopic images of Giemsa-stained human blood smears infected with Plasmodium vivax, containing approximately 80,000 cells. The LIVECell dataset includes 5,239 expert-annotated phase-contrast microscopy images, comprising 1,686,352 annotated cells across eight cell types, with predefined splits for training, validation, and testing, along with various subset sizes for scalability experiments.
Noteably that both the BloodCell and LIVECell datasets present segmentation tasks that are relatively simpler compared to our primary dataset due to the consistent cellular morphologies observed. Despite these variations, our MorphoSeg method achieves promiss performance improvements, as summarized in Table 4, highlighting its effectiveness across diverse cellular imaging conditions.